OpenStudy (anonymous):

Which mathematical sentence represents the solution for d in the equation 6e = ef + 3d?

OpenStudy (anonymous):

6e = ef + 3d 6e - ef = 3d (6e - ef)/3 = d

OpenStudy (ajprincess):

e(6-f)/3=d

OpenStudy (anonymous):

thanks guys!